Database last updated on Sun, 01 Feb 2026 12:10:51 +0000 by localhost

Instrument constant changes for the week before 2023314 (2 sec, 0 rows)

HG burned out? (AP results when HG is ON) (0 sec, 17 rows)

Measurements and tests summary.


_jday_loc_serial_skcaodedsdtedfzhgrssluvzs
314STONY PLAIN#055ecsc,Next schedule will be ecsc (now ecsc ZS= 93 ZS= 93.175)3838121020164617
(0 sec)

Dead time (1 sec)

File nameCountSerialExpected (+/-5) Measured 30-day history
B31423.0551#0553642